# GPTKB: Building Very Large Knowledge Bases from Language Models

### v2

@@ Abstract
General-domain knowledge bases (KB), in particular the “big three” –Wikidata, Yago and DBpedia– are the backbone of many intelligent applications. While these three have seen steady development, comprehensive KB construction at large has seen few fresh attempts.
In this work, we propose to build a large general-domain KB entirely from a large language model (LLM). We demonstrate the feasibility of large-scale KB construction from LLMs, while highlighting specific challenges arising around entity recognition, entity and property canonicalization, and taxonomy construction. As a prototype, we use GPT-4o-mini to construct GPTKB, which contains 105 million triples for more than 2.9 million entities, at a cost 100x less than previous KB construction projects.
Our work is a landmark for two fields: For NLP, for the first time, it provides constructive insights into the knowledge (or beliefs) of LLMs. For the Semantic Web, it shows novel ways forward for the long-standing challenge of general-domain KB construction. GPTKB is accessible at http://gptkb.org.
GPTKB: Building Very Large Knowledge Bases from Language Models
Yujia Hu1   Shrestha Ghosh2   Tuan-Phong Nguyen2   Simon Razniewski1 1ScaDS.AI & TU Dresden, Germany 2Max Planck Institute for Informatics, Saarbrücken, Germany yujia.hu@mailbox.tu-dresden.de  {ghoshs,tuanphong}@mpi-inf.mpg.de  simon.razniewski@tu-dresden.de
@@ 1 Introduction
Figure 1: Overview of our approach for LM-based KB construction.
General-world knowledge bases (KB) like Wikidata Vrandecic and Krötzsch (2014), Yago Suchanek et al. (2007) and DBpedia Auer et al. (2007) are important backbones for intelligent applications. While these projects exist for over a decade, innovation in the field of general-world KB construction is low, with neither fundamental paradigm shifts nor major novel projects emerging Weikum et al. (2021).
Recently, large language models (LLMs) stirred up many fields of AI Bubeck et al. (2023), and also been proposed as sources for structured knowledge Petroni et al. (2019). More specifically, Cohen et al. (2023) showed how, in principle, one can build a KB from an LLM by simple factual prompts and iterative graph expansion, but have not attempted this at scale.
The success of LLMs has also raised important intrinsic questions, in particular, what and how much these models know or believe Petroni et al. (2019); Jiang et al. (2020); Roberts et al. (2020); Veseli et al. (2023); Sun et al. (2024); Wu et al. (2024b). A large set of benchmarks and studies investigate this via example-based prompting, e.g., to determine, how many answers to common benchmarking question answering (QA) datasets are known by LLMs. However, all these works remain non-exhaustive, investigating samples from specific datasets or domains, without attempting to materialize all knowledge of an LLM.
In this paper we propose to comprehensively materialize the knowledge/beliefs [^1] The terminology here is contentious, see Section 6.1. of LLMs into a KB. In particular, we propose to use iterative graph expansion to obtain the complete set of (named-entity-centric) LLM knowledge, and to consolidate that knowledge with LLM-based entity disambiguation, class and relation canonicalization, and taxonomy induction (see Fig. 1). This proposal faces several challenges:
@@
    Termination/cost/runtime: State-of-the-art KBs contain millions of entries, and LLMs are known to hallucinate. It is therefore not clear if and when iterative graph exploration will terminate, and how to perform it under practical monetary and time constraints.
@@
    Flexible and relevant knowledge elicitation: LLMs possess a wide variety of knowledge, and the amount of knowlege per entity varies. We need a method that elicits as much general-world knowledge as possible, without encouraging hallucinations, and without falling into bottomless corners, e.g., open-ended phrases or translations.
@@
    Canonicalization and disambiguation: Hallmarks of existing KBs are disambiguated entities and coherent classes, relations and taxonomies. Iterative prompting risks surfacing expressions that are not globally coherent.
Our approach builds on the following ideas: To overcome scaling issues, and obtain relevant knowledge, we utilize a commercial API that allows to massively send batch requests (GPT-4o-mini), and utilize named entity recognition (NER) and carefully crafted prompts to restrict the explored spaces, along with prompts that allow varied answer sizes. To obtain a coherent KB, we perform a set of canonicalization and disambiguation steps, entirely relying on the LLM itself. In summary, our salient contributions are:
@@
    To the best of our knowledge, we are the first to propose to construct comprehensive KBs entirely from LLMs.
@@
    We develop a KB construction pipeline that overcomes termination, quality and runtime issues.
@@
    We construct GPTKB, the first large-scale KB entirely built from an LLM, containing over 104M assertions for over 2.9M entities.
Our work is a significant step for two communities: For the NLP community, for the first time, we provide a proof-of-concept methodology that enables constructive insights into what LLMs know (or believe). For the Semantic Web community, we provide fresh momentum for the long stale task of open-domain KB construction, and provide code, a concrete resource, GPTKB, both as a 3.8 GB download, and via an online browsing interface and SPARQL query interface at http://gptkb.org.
@@ 2 Related work
Large-scale KB construction
@@
GPTKB 2.9M 105M
Table 1: Size comparison1 of major KBs. The ones in the top block are Wikimedia-related, NELL and ReVerb are text-extracted.
11footnotetext: Sources: Wikidata: See https://www.wikidata.org/wiki/Wikidata:Statistics and https://grafana.wikimedia.org/d/000000175/wikidata-datamodel-statements. Wikidata5m: Wang et al. (2021). Yago4.5: Suchanek et al. (2024). DBpedia: English version as per Table 2 in Lehmann et al. (2015). NELL: As per Figure 5 (left) in Mitchell et al. (2018). Reverb: As per Lin et al. (2012) and http://web.archive.org/web/20220307185343/https://openie.allenai.org/
Extent of LLM knowledge
Since the emergence of LLMs, the question of how much these models know has frequently been raised Roberts et al. (2020); Jiang et al. (2020); Veseli et al. (2023); Sun et al. (2024); Wu et al. (2024b). So far, the widely adopted approach is to sample a domain of interest, e.g., question-answer pairs from a common benchmark, or triples from Wikidata, probe the LLM, and compute the fraction that the LLM can correctly answer/complete. Yet this does not provide an overall estimate of LLM knowledge, as the LLM might hold a lot of knowledge outside the focus of existing benchmarks. For instance, we found that GPT-4o-mini holds substantial knowledge about people’s hobbies, which are not well covered in existing KBs.
Knowledge extraction from LLMs
Most previous works only transiently probe LLMs for individual objects of subject-relation-pairs like Vannevar_Bush-citizenship Petroni et al. (2019); Veseli et al. (2023); Sun et al. (2024); Wu et al. (2024b), and use samples to estimate the LLMs knowledge. Funk et al. (2023) explore ab-inito taxonomy construction from LLMs. Kassner et al. (2021) propose a small-scale persistent memory component for ensuring LLM answers remain consistent over multiple prompts. Nguyen et al. (2024) iteratively prompt GPT-3.5 for obtaining 167K sentences containing cultural commonsense for 11K subjects. Cohen et al. (2023) propose to iteratively prompt GPT-3 for relations and relational assertions for triples. In difference to Nguyen et al. (2024), we are after structured content (triples), at a much larger scale. In difference to Cohen et al. (2023), we unveil the practical challenges that such a graph exploration faces, tackle them, and perform LLM-based KB construction at scale.
@@ 3 Methodology
An overview of our approach is shown in Fig. 1. In the first phase, we iteratively elicit LLM triples for a given subject, and enqueue newly encountered named entities for further triple elicitation. In the second phase, we consolidate the resulting triple set by canonicalizing entities, relations and classes, and by constructing a coherent taxonomy. Our paradigm is that as much knowledge as possible should come from the LLM itself, we therefore refrain from imposing any standardized vocabularies, or using existing KBs as disambiguation references.
@@ 3.1 Knowledge elicitation
We start the extraction process from a single seed subject, Vannevar Bush, the visionary behind the concept of hyperlink-based knowledge organization Bush (1945). From the triples obtained for him, just like in web crawling, we can then identify further entities (e.g., MIT (affiliation) or Everett, MA (birth place)), for which we can then elicit further knowledge, and so on Cohen et al. (2023). Figure 2 illustrates how in 3 hops, we arrive at entities of diverse types, such as historical event (Boston Tea Party, Manhattan Project), newspaper (The Times) and magazine (The New Republic).
Figure 2: Graph exploration from seed entity Vannevar Bush.
Knowledge prompting
A major challenge in knowledge elicitation is to elicit as much knowledge as possible, but at the same time, not encouraging hallucinations. We found that without guidelines on the expected number of triples, LLMs systematically returned much fewer triples than they could correctly return. However, a static number would not be suitable either, as e.g., for Albert Einstein, an LLM should return many more triples than for the average person. We solve this via a subtask where the LLM first specifies the number of triples it believes to know for the subject, with two ranges as orientation. In difference to Cohen et al. (2023), we also streamlined the process by dropping the separate elicitation of relations and relation-specific triples. To structure the knowledge, we also request that at least one instanceOf triple should be returned. Output parsing is eased by using the structured outputs feature of OpenAI’s API. In turn, this also reduces hallucinations, e.g., additional qualifiers, textual descriptions, or similar. The full prompt is presented in Fig. 3.
Figure 3: Our prompt for knowledge elicitation.
@@
Our early attempts at graph exploration were plagued by topical runaway into linguistic knowledge, translations, etc. To keep the process on target, we use spaCy’s NER module to enqueue only objects who are among common classes of named entities. We include language-specific NER models for 10 major languages, and employ a lightweight rule-based filter to exclude websites URLs and other noise. The NER component appears to still hold room for improvement, likely due to the fact that NER models are traditionally trained on sentences, while in our task, the inputs are only the short, context-free entity labels.
@@ 3.2 Knowledge consolidation
Because input prompt size directly relates to the complexity and cost of prompting, it is not possible to feed a large amount of existing entity or taxonomy knowledge into the knowledge elicitation phase. Consequently, the output exhibits a huge degree of redundancy and variance, which we consolidate post-hoc.
@@
The naive process generates 688,855 distinct relation names, with many obvious duplicates, e.g., instanceOf, isA, or InstanceOf. To remove redundancies within this set, we apply greedy clustering, based on embedding similarity computed via SentenceTransformers Reimers and Gurevych (2019). Specifically, we compare each relation $R$ with all more frequent ones, starting from the most frequent one, and merge relation $R$ into relation $S$, if the textual embedding similarity between the names of $R$ and $S$ is greater than:
$1.2\times\frac{\log{(\textit{freq}(R))}}{\log{(\textit{freq(most\_freq\_relation)})}}.$
In particular, the formula is designed so that the similarity threshold varies with the frequency of the relation, leading to a more aggressive removal of relations that occur with low frequency. This way, we reduce the number of relations to 2,133 (a 323x reduction).
@@
The naive process also generates 83,770 distinct class names (objects for instanceOf), with both obvious duplicates, and many overly specific cases. We employ the same greedy clustering as for relations, with the same multiplier 1.2 in the merging criterion. This way, we reduce the number of classes to 367.
@@
The classes in the KB so far do not form a coherent, or even connected taxonomy, as the knowledge elicitation process only expands named entities. But we also cannot simply prompt for a complete taxonomy to put on top (e.g., following the approach in Funk et al. (2023)), because we need to include the existing classes. We chose to proceed in three steps: In step 1, we prompt the LLM to generate an initial high-level taxonomy of about 50 classes (see prompt in Fig. 4 (top)). In step 2, we prompt the LLM to assign to each class in the KB a score describing how general it is (see prompt in Fig. 4 (upper middle)), then use that score to rank the classes from most general to most specific. In step 3, we insert classes into the taxonomy, one at a time. We initially tried this with the full taxonomy as input, but found that the LLM struggled. We therefore now proceed recursively. Starting from the nodes below the root, we ask the LLM repeatedly, whether the given class to insert belongs below one of them, in which case we repeat this process recursively with that node’s children (see prompt in Fig. 4 (lower middle)). The recursion stops either when the given class is not a subclass of any child anymore (in which case it is inserted as sibling), or when a leaf is reached (in which case it is inserted as new child). Then we prompt LLM to update the (sub)taxonomy from current node with the given class (see prompt in Fig. 4 (bottom)). We feed the final taxonomy back into GPTKB via the rdfs:subClassOf relation. Further LLM-based refinement in the style of Peng et al. (2024) could be considered.
You are a knowledge base construction expert. Your task is to initialize a seed taxonomy with general categories, which you will update later with given classes. Please return only the seed taxonomy in json form with indentation: Class: <new class> You are a knowledge base construction expert. Your task is to create a taxonomy for a knowledge base. Beforehand, you need to give each given class a score describing how general it is. The score is a integer ranging only from 1, for the most general concept, to 10, for the most specific concept. Please return only the score of the given class: Candidate branches: <branches of current node> Class: <class to add> You are a knowledge base construction expert. Your task is to integrate a given class into the taxonomy. If the given class is a subclass of one of the candidate branches, return only the exact name of that branch. Otherwise, return only NULL. Taxonomy: <taxonomy from current node> Class: <class to add> You are a knowledge base construction expert. Your task is to update the given taxonomy with the given class. You can consider the categorization of the taxonomy, but you can not modify the names of the classes in the taxonomy. Please return only the updated taxonomy in json form:
Figure 4: Prompts for seed taxonomy construction (top), class generability scoring (upper middle), recursive class insertion check (lower middle), and (sub)taxonomy update (bottom).
@@
Naive graph exploration frequently yields duplicates, for instance, John_F._Kennedy and John_Fitzgerald_Kennedy. To remove these duplicates without requiring extraordinary runtime, we follow the standard blocking-based deduplication approach Köpcke and Rahm (2010). This technique is usually based on choosing an appropriate blocking key: one attribute by whose values we obtain meaningful partitions on entities. The key should meet two criteria: on the one hand, it should be a necessary condition for equivalence; on the other hand, it should be distinctive enough to partition the entities into not too large blocks.
We focus on entities that are instances of the most interesting class, humans. We block entities by birth dates. [^2] Only considering those with a birth date at a resolution more specific than a year. Within each block, we consider an entity pair as duplicates if both their labels exhibit at least 0.85 similarity using SentenceTransformers, and 30% of their triples are exactly the same. We then merge these entities, keeping the name of the entity that was created first. In total, this step removes 6,987 duplicate instances of Person. More advanced methods could utilize LLMs themselves for deciding on entity equivalence Ding et al. (2024).

### v3

# GPTKB: Comprehensively Materializing Factual LLM Knowledge

@@ Abstract
LLMs have majorly advanced NLP and AI, and next to their ability to perform a wide range of procedural tasks, a major success factor is their internalized factual knowledge. Since Petroni et al. (2019), analyzing this knowledge has gained attention. However, most approaches investigate one question at a time via modest-sized pre-defined samples, introducing an “availability bias” Tversky and Kahneman (1973) that prevents the discovery of knowledge (or beliefs) of LLMs beyond the experimenter’s predisposition.
To address this challenge, we propose a novel methodology to comprehensively materialize an LLM’s factual knowledge through recursive querying and result consolidation.
As a prototype, we employ GPT-4o-mini to construct GPTKB, a large-scale knowledge base (KB) comprising 101 million triples for over 2.9 million entities - achieved at 1% of the cost of previous KB projects. This work marks a milestone in two areas: For LLM research, for the first time, it provides constructive insights into the scope and structure of LLMs’ knowledge (or beliefs). For KB construction, it pioneers new pathways for the long-standing challenge of general-domain KB construction. GPTKB is accessible at https://gptkb.org.
GPTKB: Comprehensively Materializing Factual LLM Knowledge
Yujia Hu1   Tuan-Phong Nguyen2   Shrestha Ghosh3   Simon Razniewski1 1ScaDS.AI & TU Dresden, Germany 2Max Planck Institute for Informatics, Saarbrücken, Germany 3 University of Tübingen, Germany yujia.hu@mailbox.tu-dresden.de  {tuanphong,ghoshs}@mpi-inf.mpg.de  simon.razniewski@tu-dresden.de
@@ 1 Introduction
{NiceTabular}
p0.44|p0.48 Existing LLM analyses Our proposal
Prompt-and-discard:
- Transient single-focus explorations (e.g., factuality OR bias OR cutoff) Persistent resource:
- Materialized KB reusable for a wide array of questions
Sample availability bias:
- Insights bound to experimenter’s predisposition Recursive materialization:
- Discover unconceived LLM knowledge/beliefs
Scratching the surface:
- Few 100s-100K samples
- Breadth and depth of LLM knowledge untouched Massive-scale:
- Over 100M records
- Recursive crawl to unprecedented breadth and depth
Table 1: Comparison of existing LLM knowledge analysis approaches and our proposal.
LLMs have been one of the most exciting recent developments in NLP and AI, and next to their ability to perform a wide set of procedural tasks, a major success factor is the factual knowledge that they have internalized Bubeck et al. (2023). Their potential to store factual knowledge, like (Nelson Mandela, award, Nobel Peace Prize), was first highlighted by Petroni et al. (2019), and this has generated an own (sub-)field of studying all aspects of factual knowledge in LLMs, from trying to locate it, to estimates of their storage potential, to techniques to effectively elicit it Jiang et al. (2020); Roberts et al. (2020); Veseli et al. (2023); Sun et al. (2024); Wu et al. (2024). A large set of benchmarks and studies transiently investigate the knowledge storage ability via example-based prompting, e.g., to determine, how many triples from common knowledge bases (KBs) or question answering (QA) datasets are known to LLMs. However, all these works are subject to an availability bias Tversky and Kahneman (1973), i.e., they only surface LLM knowledge/beliefs [^1] The terminology here is contentious, see Section 6.1. on topics that were known to and pre-designed by the experimenter. They inherently cannot discover knowledge on topics outside of their benchmarks.
Figure 1: Overview of our approach for factual knowledge materialization from LLMs.
To enable comprehensive insights into LLM knowledge, we propose to extensively materialize the knowledge of LLMs into a KB (see Table 1). General-world KBs like Wikidata Vrandecic and Krötzsch (2014), Yago Suchanek et al. (2007) and DBpedia Auer et al. (2007) are important and long-standing backbones for AI applications, yet have seen little innovation in the last years, and have only been constructed manually, or by scraping semistructured web content Weikum et al. (2021). In particular, we propose to use recursive graph exploration to obtain comprehensive (named-entity-centric) LLM knowledge, and to consolidate it via LLM-based entity disambiguation, class and relation canonicalization, and taxonomy induction (see Fig. 1). This proposal faces several challenges:
@@
    Runtime and cost: State-of-the-art KBs contain millions of entries, and LLM inference is relatively slow. It is therefore unclear how to perform comprehensive knowledge elicitation under practical monetary and time constraints.
@@
    Variance, hallucinations, and scoping: Latent knowledge in LLMs covers a wide set of topics, varies by entity, and includes hallucinations. We need a method that elicits a high quantity, without encouraging hallucinations, and without falling into bottomless corners, e.g., open-ended phrases or translations.
@@
    Global inconsistency: Distinctive prompts risks surfacing outputs that are not globally consistent, e.g., introducing duplicate relations, entities, or disconnected class hierarchies.
Our approach builds on the following ideas: To overcome scaling issues, and obtain relevant knowledge, we utilize a commercial API that allows to massively batch requests, and utilize named entity recognition (NER) and carefully crafted prompts to restrict the explored space, along with prompts that encourage varied answer sizes. To obtain a coherent KB, we perform a set of canonicalization and disambiguation steps, entirely relying on the LLM itself. In summary, our salient contributions are:
@@
    To the best of our knowledge, we are the first to propose extensive materialization for analyzing factual LLM knowledge.
@@
    We present a scalable methodology to elicit massive amounts of LLM knowledge, and to consolidate it.
@@
    Using GPT-4o-mini, we construct GPTKB, the first large-scale KB entirely built from an LLM’s parametric knowledge, containing 101M assertions for over 2.9M entities.
  4. 4.
    We use GPTKB to exemplarily analyze GPT-4o-mini’s factual knowledge in terms of scale, accuracy, bias, and cutoff, at the same time.
Our work is a significant advancement for two areas: For LLM research, we introduce a proof-of-concept methodology that, for the first time, enables constructive insights into the knowledge (or beliefs) of LLMs. For the Semantic Web, we provide fresh momentum for the long stale problem of open-domain KB construction. We provide code and a concrete resource, GPTKB, both as a 3.8 GB download, and via an online browsing interface and SPARQL query interface at https://gptkb.org.
@@ 2 Related work
Extent of LLM knowledge
Since the emergence of LLMs, the question of how much these models know has frequently been raised Petroni et al. (2019); Roberts et al. (2020); Jiang et al. (2020); Veseli et al. (2023); Sun et al. (2024); Wu et al. (2024). So far, the widely adopted approach is to sample a domain of interest, e.g., question-answer pairs from a common benchmark, or triples from Wikidata, transiently probe the LLM, and compute the fraction that the LLM can correctly answer/complete. This is prone to “availability bias” Tversky and Kahneman (1973) since we do not get a comprehensive view of the LLM knowledge outside the focus of existing benchmarks. For instance, we found that GPT-4o-mini holds substantial knowledge about art periods or hobbies, which are not covered in existing KBs. Kassner et al. (2021) propose a small-scale persistent memory component for ensuring LLM answers remain consistent over multiple prompts. Nguyen et al. (2024) iteratively prompt GPT-3.5 for obtaining 167K sentences containing cultural commonsense for 11K subjects. Cohen et al. (2023) propose to iteratively prompt GPT-3 for relations and relational assertions for triples. In difference to Nguyen et al. (2024), we are after structured content (triples), at a much larger scale. In difference to Cohen et al. (2023), we unveil the practical challenges that such a graph exploration faces, tackle them, and perform LLM-based KB construction at scale.
Iterative information extraction
Iterative information extraction is a long-standing idea, already prototyped by Google cofounder Sergey Brin via the DIPRE system Brin (1998). The Snowball system by Agichtein and Gravano (2000) presented a large-scale implementation of this idea, harvesting tuples from over 300,000 newspaper documents. The KnowItAll system Etzioni et al. (2004) followed a similar approach, and led to the ReVerb KB Fader et al. (2011). Despite a proposal by Cohen et al. (2023), to date, we are not aware of any attempts to extract knowledge from LLMs at scale.
@@
Generative
GPTKB 2.9M 101M
Table 2: Size comparison of major KBs. See Appendix F for sources.
Large-scale KB construction
@@ 3 Methodology
An overview of our approach is shown in Figure 1. In the first phase, knowledge elicitation, we iteratively elicit LLM triples for a given subject, and enqueue newly encountered named entities for further triple elicitation. In the second phase, knowledge consolidation, we consolidate the resulting triple set by canonicalizing entities, relations and classes, and by constructing a coherent taxonomy. Our paradigm refrains from imposing any standardized vocabularies, or using existing KBs as disambiguation references to effectively materialize the parameterized LLM knowledge.
@@ 3.1 Knowledge elicitation
The process starts from one or a set of seed subjects, e.g., Vannevar Bush, the visionary behind the concept of hyperlink-based knowledge organization Bush (1945). From the objects in the triples obtained for him, we can then identify further entities (e.g., MIT (affiliation) or Everett, MA (birth place)), for which we can then elicit further knowledge, and so on Brin (1998). Figure 2 illustrates how in 3 hops, we arrive at entities of diverse types, such as historical event (Boston Tea Party, Manhattan Project), newspaper (The Times), and magazine (The New Republic).
Figure 2: Graph exploration from the seed entity Vannevar Bush.
Knowledge elicitation
A major challenge is to elicit as much knowledge as possible, but at the same time, not encourage hallucinations. We found that without guidelines on the expected number of triples, LLMs showed too little variance in the number of triples per subject, while we would expect them to return many more triples for Einstein than for other entities. We solve this via indications that are defined in relation to the entity’s popularity. In difference to the proposal Cohen et al. (2023), we also drop the separate relation elicitation, and relation-specific knowledge extraction, as these do not scale. To structure the knowledge, we also request that at least one instanceOf triple should be returned. Output parsing is eased by using the structured outputs feature of OpenAI’s API. In turn, this also reduces hallucinations, e.g., additional qualifiers, textual descriptions, or similar. The full prompt is in Figure 8 in the Appendix.
@@
Our early attempts at graph exploration were plagued by topical runaway into linguistic knowledge, translations, etc., because LLM-generated objects cover a wide range of imprecisely delineated data types. We experimented with various ways to filter non-named entities from the expansion, but found that existing NER frameworks struggled with the context-free entity labels available from our crawl process (most NER models are trained on sentences). In the end, we used the LLM itself to identify named entities, processing multiple candidates at once. The full prompt is presented in Figure 9 in the Appendix.
@@ 3.2 Knowledge consolidation
The knowledge elicitation phase returns a huge degree of redundancy and variance. Feeding a large amount of existing entities or taxonomy into the knowledge elicitation prompt could mitigate this problem. However, that would also increase costs significantly, and even be infeasible once we have generated millions of entities and triples. Instead, we introduce several steps to consolidate LLMs output post-hoc.
@@
The elicitation phase generates 788K distinct relation names, with many obvious duplicates, e.g., instanceOf, isA, or InstanceOf. We apply a greedy clustering algorithm to this set (see Algorithm 1). This process merges relation $r$ into a more frequent relation $s$, selected as the one with the highest textual embedding similarity to $r$ among all relations more frequent than $r$, if the similarity is greater than an adaptive threshold (defined in line 6 in Algorithm 1). This threshold varies with the frequency of the relation, leading to a more aggressive removal of relations with low frequency.
Algorithm 1 Relation clustering
1:A set of relations $\mathcal{R}$, hyperparameter $\alpha$
2:A map from relation to cluster-id $\mathcal{C}$
[…]
4:$\mathcal{C}\leftarrow\{\}$
[…]
6:for $r\in\mathcal{R}$ do
7:     $s^{\prime}=\text{argmax}([\text{similarity}(r,s)\textbf{ for }s\in\mathcal{R}\textbf{ upto }r])$
8:     $\textit{threshold}\leftarrow\alpha\times\frac{\log(\text{frequency}(r))}{\log(\text{frequency}(\text{first}(\mathcal{R})))}$
9:     if $\text{similarity}(r,s^{\prime})>\textit{threshold}$ then
10:         $\mathcal{C}[r]\leftarrow\mathcal{C}[s^{\prime}]$
11:     else
[…]
14:     end if
15:end for
16:return $\mathcal{C}$
@@
Similarly to relations, there are 103K distinct class names (i.e., objects for instanceOf relations) generated by the knowledge elicitation phase, with both obvious duplicates, and many overly specific cases. Algorithm 1 is also applied to clean this set.
In our experiment, both relation and class clustering uses SentenceTransformers Reimers and Gurevych (2019) for embedding cosine similarity computation, with $\alpha=1.2$ chosen via manual inspection of small held-out sets.
@@
The classes in the KB so far do not form a coherent, or even connected taxonomy, as the knowledge elicitation process only expands named entities. We propose Algorithm 2, which is based on the LLM itself, to build a complete taxonomy from these individual classes.
Algorithm 2 Taxonomy construction
1:Knowledge base (KB)
2:Taxonomy
3:function construct_taxonomy(KB)
4:     taxonomy $\leftarrow$ LLM_get_high_level_taxonomy()
5:     for class $\in$ KB.classes do
6:         class.generality_score $\leftarrow$ LLM_get_score(class)
7:     end for
8:     classes $\leftarrow$ sort(KB.classes, key=generality_score)
9:     for class $\in$ classes do
10:         insert_class_recursive(taxonomy.root, class)
11:     end for
12:     return taxonomy
13:end function
14:
15:function insert_class_recursive(node, class)
16:     if has_children(nodethen
17:         next $\leftarrow$ LLM_superclass(class, node.children)
18:         if next is not null then
19:              insert_class_recursive(next, class)
20:         else
21:              LLM_update_taxonomy(node, class)
22:         end if
23:     else
24:         LLM_update_taxonomy(node, class)
25:     end if
26:end function
The construction starts by generating a high-level taxonomy via the LLM. Then, for each of the existing classes in the KB (sorted by generality scores given by the LLM), we find the lowest node of which the given class is identified as a subclass, via depth-first search (see function insert_class_recursive in Algorithm 2). Then, we ask the LLM to update the sub-taxonomy starting from the found node with the given class. In this step, the LLM may generate new intermediate classes on the path from the found node to the given class.
The prompts used for seed taxonomy generation, generality scoring, superclass identification, and taxonomy update are provided in Figure 10 in Appendix C. Further LLM-based refinement in the style of Peng et al. (2024) could be considered.
@@
Naive graph exploration frequently yields duplicates, such as John F. Kennedy and John Fitzgerald Kennedy. To remove these duplicates without requiring extraordinary runtime, we follow the standard blocking-based deduplication approach by Köpcke and Rahm (2010). This approach is based on choosing a blocking key, which is an attribute by whose values we obtain meaningful partitions on entities.
We focus on entities that are instances of the most interesting class, humans. We choose to block entities by birth dates. Within each block, we consider an entity pair as duplicates if (i) their labels have close meanings (i.e., cosine similarity between SentenceTransformers embeddings are greater than 0.85), and (ii) 30% of their triples are exactly the same. More advanced methods could utilize LLMs themselves for deciding on entity equivalence Ding et al. (2024), but it would incur more cost and longer runtime.